OpenStudy (anonymous):

Use implicit differentiation to find the slope of the tangent line to the curve

OpenStudy (anonymous):

\[\frac{ y }{ x-8y }= x^2- 7\]

OpenStudy (anonymous):

im not sure if should do quotient rule or if i can multiply the denominator to the other side BY THE WAY THE CORDINATE IS (\[\left( 1,\frac{ -6 }{ -47 } \right)\]

OpenStudy (anonymous):

I would multiply to the other side first to make it easier to preform

OpenStudy (anonymous):

Yea just use quotient rule on the one side

OpenStudy (anonymous):

The equation to find the slope is dy/dx = (2x^2-8yx+y)/(x-16y) I think

OpenStudy (anonymous):

you can simplify it further but I got things to do lol

OpenStudy (anonymous):

ok ill see, thanks though

OpenStudy (anonymous):

No problem